The largest angle of a triangle is 40° larger than the second angle, which is three times the size of the smallest angle. What are the sizes of the three angles of this triangle?
Answer by rfer(16322)

1st=x
2nd=3x
3rd=3x+40
-----
x+3x+3x+40=180
7x=180-40
x=140/7
x=20
3x=60
3x+40=100